OpenStudy (anonymous):

5x+52=62 WHAT IS THE VALUE OF X

OpenStudy (solomonzelman):

subtract 52 from both sides nd divide both sides by five

OpenStudy (anonymous):

i don't get it.

OpenStudy (solomonzelman):

5x+52-52=62-52

jigglypuff314 (jigglypuff314):

do what solomon told you 5x+52=62 subtract 52 from both sides -52 -52 5x = 10 divide both sides by five /5 /5 x = 10/5 = 2

OpenStudy (solomonzelman):

get it now?

OpenStudy (anonymous):

yes.

OpenStudy (solomonzelman):

can you do another example based on this one?

OpenStudy (squirrels):

Ya so basically your trying to "x" by itself. To do that you subtract 52 by itself, but in order to do that you must subtract it from the other side as well. 5x+52=62 -52=-52 5x=10 Now to get x by itself this time divide "5x" (since 5 is multiplying x, and the opposite of multiply is divide) by 5. and like last time, what ever you do to one sides you must do to both. 5x=10 ------ 5 x=2 I can't explain any better than that :p